One number is 10/3 larger than the other. The doubled sum of these numbers is 19 1/2 . What is the smallest number?
Answer by addingup(3677)

y = 10/3x
2(x+y) = 19 1/2
2(x+10/3x) = 19 1/2
2x+20/3x = 19 1/2
6x+20x = 58 1/2
26x = 58 1/2
x = 2 1/4
and
y = 10/3*2 1/4 = 7 1/2
--------------------------------\
Check:
The doubled sum of these numbers is 19 1/2
2(2 1/4+7 1/2) = 19 1/2 Correct
:
John